Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
nardecchia, rocchi - 20:05 Monday 07 October 2019 (47135) Print this report
CO2 lasers activities

The power calibration cannot be performed today because after the turning off of the Cinogy (# 47120) a long thermal transient started. So we decided to start restoring the CH line in view of the increase of input power foreseen at the end of this week.

List of the activities on the benches:
We installed the lens of f=140mm in order to have the beam size on the CP equal to 50 mm (the same beam size of the YAG on the ITMs).
A rough centering has been done with the thermal camera by taking as reference the DAS.
We also installed the pre-amplifier on the beam sampler at the output of the laser (TCS_NI_CO2_PWRLAS).
The new threshold in the DMS has been updated accordingly.

We cleaned the bench from the biological rests of a dead mouse (see attached figure) and we removed the 'old' lens on the CH line (f=50 mm in order to have the beam size on the CP equal to 150mm).

1)power calibration.
1)installation of the nominal lens (f=140mm in order to have a beam size on the CP equal to 50 mm).
2)installation of the same pre-amplifier already installed on the NI bench at the output of the WI laser.
3)power calibration

We will continue tomorrow morning during the maintenance.


Images attached to this report
Comments to this report:
nardecchia, rocchi - 18:50 Tuesday 08 October 2019 (47150) Print this report

This morning we performed the power calibration on both DASes after the end of the thermal transient caused by the switching off of the infrared cameras.

We left the DAS powers projected on the CPs values optimized for the input power of about 19 W:

NI--> INNER: 225mW

         OUTER:2 W

WI--> INNER: 100 mW

         OUTER: 68 mW

nardecchia, rocchi - 0:00 Friday 11 October 2019 (47184) Print this report

WI CO2 CH alignment

This morning we installed the lens needed to obtain the same beam size of the YAG on the WI bench and roughly aligned the central heating by using the DAS as reference.

To fine tune the CH alignment we switched on the WI RH at 10.15 UTC by applying a power of 20 W to evaluate center of the mirror in the HWS-INJ maps.

Then we turned on the CH and it seems well aligned on the CP centre.

NI CO2 CH alignment

We turned on the NI RH at 13.05 UTC by applying  20W.

We noticed that the OPL due to the RH was quite off-centered  so we decided to re-align the HWS-DET SLED  on the TM by moving the mirror on EDB.

To do this, we used the CH since it is faster wrt the RH  (fig.2).

Then we turned off the RH (15.37 UTC) and CH in order to cool down TM+CP.

At 17.28 UTC we turned on again the RH and we compared the OPL map of the RH with the CH one (fig. 3) and they seems quite well aligned.

As planned we left the TCS completely off.



Images attached to this comment
mantovani, bersanetti, nardecchia - 10:44 Friday 11 October 2019 (47189) Print this report

In order to perform these measurement the SDB1 drift control on B5 QPD has been disabled manually (the LC set points have been updated as the on-line ones). Then the enabling of the drift control has been disabled in to let TCS people to work on CH even if the ITF unlocks (re-enable rigth now).

Search Help